# MS-0000-0000.00

### v1

@@ 4.3.4   Discussion Extension: Considering Team and Task Complexities
While our discussion of future research has primarily focused on a simpler context, the increasing complexity of team dynamics due to variations in size and composition warrants further brief elucidation. It is essential to consider how future studies can be extended to address the challenges posed by complex team and task configurations. Specifically, smaller teams with fewer members and one or two AI agents often benefit from streamlined communication and faster decision-making, making them effective for tasks requiring agility and close collaboration. Conversely, larger teams that bring a broader range of expertise face challenges such as coordination difficulties, slower decision-making, and increased risks of misalignment. This highlights the need for enhanced coordination mechanisms and adaptive workflows to maintain alignment and cohesion in larger teams. Moreover, teams involving multiple humans and multiple AI agents introduce dynamic interactions that heighten creativity but also require advanced strategies to manage diverse inputs and prevent cognitive overload. In addition to team size, task complexity significantly influences team formulation and coordination. Tasks with interdependencies across various objectives demand structured workflows that can dynamically adjust to AI-generated insights. For instance, AI contributions in one task area may ripple into changes in related areas, requiring seamless transitions and alignment across team stages. Addressing these complexities underscores the critical role of role specification in managing interdependencies and ensuring alignment with overarching goals.
@@
Considering these complexities, building on the discussions in Sections 4.3.2 and 4.3.3, future research should further (1) explore the optimal configurations of team size and composition for different task complexities, examining how these factors impact team dynamics and effectiveness (Wu et al. 2019b, Yanai and Lercher 2024), and (2) investigate mechanisms for managing task interdependencies and role transitions in dynamic, multi-functional human-AI teams. These efforts will provide insights into addressing the increasing complexity of modern human-AI teams while optimizing specification and coordination.
4.4 Implications on Social Dynamics in Teams
@@
Conflicts are inherent in all teams, whether human-only or human-AI. These conflicts typically arise from task, relational, and process-related dynamics (de Wit et al. 2013). In human-AI teams, such conflicts will continue to affect performance, but the generative or causal factors require new theoretical insights. Human-AI teams present new nuances, such as increased role fluidity—the ability to adapt and transition roles based on context and task requirements. Role fluidity can influence process clarity or conflict depending on how teams implement and manage their roles. If not managed early on, process conflict contributes to relational conflict. As AI agents take on more tasks, it is reasonable to expect changes to the shared mental models among team members. To the extent the mental models cohere, reductions in task conflicts are to be expected, which in turn may reduce process conflicts. It is, however, unclear as to whether the use of AI agents in teams interrupt the formation of shared mental models among humans or if they strengthen shared mental models.
@@ 4.4.1   Expected Outcomes
Research on conflict in human-AI teams will need new theoretical lenses to study causative factors of conflict in such teams. The expected benefits of flat hierarchies and role fluidity have to be balanced against the potential harmful implications on team conflict. Extant research has raised the plausibility of increased interpersonal and status conflicts in flat organizational structures (Anicich et al. 2016). Fluidity in roles often causes process conflict, which, if not managed, leads to relational conflicts. If the key benefits of human-AI teams are to be realized, teams need to more effectively manage such conflicts. Human-AI teams ned to address imbalances in responsibility and rewards to sustain trust, engagement, and overall team effectiveness (Gabriel et al. 2024). We argue that human-AI teams need their leaders and team members to develop a stronger meta-understanding of process goals and roles as opposed to focusing on transactional level agreements on process and task execution. This, in our opinion, is the key differentiation in theorizing the social dynamics of human-AI teams.
@@ 4.4.2   Evaluations and Issues
Evaluating Conflict in Human-AI Teams:  Role fluidity enhances flexibility in task assignments and delegation but can lead to relational, process, and status-related conflicts if not well managed. Researchers should (1) explore how role fluidity leads to conflicts in human-AI teams and identify managerial actions that reduce such conflicts. For instance, AI agents can aid shared mental model formation through self-explanation and reflection, improving role clarity and minimizing conflicts.. Attitudes toward both human and AI teammates, especially when performance varies due to task allocations, are critical. For instance, SimanTov-Nachlieli (2024) emphasize that employees’ trust in AI is linked to their performance rankings, making it vital to address dynamic perceptions of self-efficacy and collaboration willingness to maintain overall team resilience. Importantly, future research should also (2) evaluate how divergent and convergent thinking styles are balanced within human-AI teams to reduce task and process conflicts, focusing on the role of AI in fostering or disrupting this balance (Lanaj et al. 2018). This includes investigating how AI systems can facilitate social learning processes by enabling team members to integrate diverse perspectives, resolve conflicts constructively, and adapt roles dynamically for cohesive decision-making (Choi et al. 2024).
@@
Information Sharing and Knowledge Management in Human-AI Teams:   Effective information sharing is fundamental to human-AI team coordination and performance, particularly for enabling role fluidity and adaptive learning. Knowledge reuse involves not only accessing archived information but also meaningfully applying it to new tasks, allowing the team to build on past experiences and insights. We draw upon Nonaka’s dynamic theory of knowledge creation to identify key research directions for human-AI teams context (Nonaka 1994). Nonaka’s model includes four knowledge interaction patterns: socialization (tacit–tacit), externalization (tacit–explicit), internalization (explicit–tacit), and combination (explicit–explicit). The inclusion of AI agents clearly has ramifications on the interaction patterns in human-AI teams as it relates to knowledge sharing and creation. Of the four patterns, perhaps the combination pattern is relatively well understood. AI agents today are highly capable of combining internal and external knowledge and synthesizing it for human consumption. Future research should (1) examine how human-AI teams can effectively organize to improve socialization, externalization, and internalization. For instance, can AI agents impede tacit knowledge formation in teams since they may interrupt or replace socialization between team members? how can organizations enhance socialization in AI-enabled teams? Additionally, (2) can the ability of AI agents to transform multimodal inputs (text, voice, video, etc.) accelerate knowledge creation through improved externalization? Similarly, (3) do interactions with AI agents during task completion impede or help internalization? Overall, human-AI teams will witness a disruption of the dyadic interactions in knowledge formation, and theoretical development in this domain is limited, making it difficult to reliably predict how organizational knowledge creation improves or deteriorates in the presence of AI agents.
@@ 4.4.3   Mechanisms and Resolutions
Fostering Adaptability and Resilience through Social Learning:  Resilient human-AI teams thrive on structured social learning processes such as reflection, experimentation, and collaborative problem-solving, which enable them to coadapt and address disruptions effectively. Team reflexivity – collective reflection on objectives, strategies, and processes – enhances problem identification and resolution, improving overall team adaptability (Schippers et al. 2014). Collaborative problem-solving fosters critical thinking and collective learning, equipping teams to navigate dynamic and uncertain environments (Van den Bossche et al. 2006). These processes are integral to dynamically aligning roles between human and AI team members, strengthening team resilience. Particularly, simulated learning environments offer a practical and scalable strategy for accelerating these social learning processes. By providing safe, controlled spaces for experimentation, these environments enable teams to engage in iterative practice, refine workflows, and test strategies without the risks of real-world tasks (Salas et al. 2009). Simulation-based learning bridges theoretical knowledge and practical application, enhancing decision-making and adaptability in complex scenarios (Wei et al. 2024). Iterative experimentation within such environments fosters mutual adaptation between humans and AI, supporting the development of shared understanding and improved collaborative efficiency. Psychological safety is a cornerstone of effective social learning. Creating environments where team members can share honest feedback and take calculated risks without fear of negative repercussions promotes openness, creativity, and innovation (Edmondson 1999). Integrating simulated learning opportunities with feedback-rich environments ensures that human and AI agents can coevolve efficiently, addressing both operational efficiency and ethical alignment (Brinkmann et al. 2023). Future research should (1) investigate how simulated learning environments can incorporate iterative experimentation – such as real-time role rotation and collaborative problem-solving exercises – to enhance role fluidity and team resilience. Additionally, (2) explore governance strategies that embed social learning processes, such as feedback loops and ethical audits, into human-AI workflows to improve adaptability, trust, and long-term alignment (Mökander and Floridi 2023).
@@ 4.4.4   Discussion Extension: Considering Goal and Task Dynamics
On top of the discussions in Sections 4.4.2 and 4.4.3, future research should (1) explore how advanced social learning frameworks can integrate conflict management and iterative feedback loops into team maintenance and training, ensuring that human-AI teams maintain alignment and cohesion as goals and tasks evolve. Additionally, (2) investigate mechanisms to enhance role fluidity by developing adaptive protocols for task transitions, interdependency management, and real-time adjustments, ensuring that both human and AI contributions remain dynamically aligned with the team’s overarching objectives. These efforts will provide actionable insights into sustaining the effectiveness, adaptability, and resilience of human-AI teams in increasingly complex and dynamic environments.
@@ 5 Concluding Remarks
The disruptive innovation of AI-based intelligent agents compels us to revisit and refine the definition of human–AI teaming. Building on an extended team SA model, we present a comprehensive framework for examining key dimensions of human-AI teaming, including formulation, coordination, maintenance, and training. We also address issues arising from team complexity, such as variations in composition, size, goals, task dynamics, and long-term interactions. Ultimately, we emphasize the importance of revisiting human-AI teaming to foster a more dynamic and interactive relationship between humans and AI agents. Our work offers a forward-looking research agenda to guide future investigations in this evolving field.
Our discussion proposes the opportunities and challenges presented by the integration of AI into collaborative settings. On the one hand, AI offers immense potential for augmenting human capabilities, improving decision-making, and fostering innovation. On the other hand, unresolved tensions, such as trust and alignment considerations, underscore the need to design systems and processes that facilitate effective and sustainable collaboration and dynamic learning in teams. As such, future research should prioritize both technical development and the socio-behavioral foundations of teaming.
@@
In conclusion, human-AI teaming represents a critical frontier in team science, offering vast opportunities for innovation and productivity while raising questions about the nature of collaboration, decision-making, and shared accountability. By addressing the gaps and tensions identified in this paper, researchers and practitioners can shape a future where human-AI teams achieve their full potential, benefiting individuals, organizations, and society at large.

### v2

@@ 4.3.4   Discussion Extension: Considering Team and Task Complexities
While our discussion has focused on simpler team structures, growing complexity in team size and composition presents new challenges for human–AI collaboration. Future research should examine how role and task allocation strategies scale in larger, more diverse teams. As complexity increases, maintaining clarity, adaptability, and alignment becomes more difficult. Investigating how task coordination performs across varying configurations can help develop robust frameworks that ensure effectiveness and coherence in increasingly dynamic environments. Specifically, smaller teams with fewer members and one or two AI agents often benefit from streamlined communication and faster decision-making, making them effective for tasks requiring agility and close collaboration. Conversely, larger teams that bring a broader range of expertise face challenges such as coordination difficulties, slower decision-making, and increased risks of misalignment. This highlights the need for enhanced coordination mechanisms and adaptive workflows to maintain alignment and cohesion in larger teams. Moreover, teams involving multiple humans and multiple AI agents introduce dynamic interactions that heighten creativity but also require advanced strategies to manage diverse inputs and prevent cognitive overload. In addition to team size, task complexity significantly influences team formulation and coordination. Tasks with interdependencies across various objectives demand structured workflows that can dynamically adjust to AI-generated insights. For example, AI contributions in one task area may ripple into changes in related areas, requiring seamless transitions and alignment across team stages. Addressing these complexities underscores the critical role of role specification in managing interdependencies and ensuring alignment with overarching goals.
@@
Considering these complexities, building on the discussions in Sections 4.3.2 and 4.3.3, future research should further (1) explore the optimal configurations of team size and composition for different task complexities, examining how these factors impact team dynamics and effectiveness (Wu et al. 2019b, Yanai and Lercher 2024), and (2) investigate mechanisms for managing task interdependencies and role transitions in dynamic, multi-functional human-AI teams. It also requires to strike a good balance between learning and productivity (Mayo et al. 2025). These efforts will provide insights into addressing the increasing complexity of modern human-AI teams while optimizing specification and coordination.
4.4 Social Learning: Implications on Social Dynamics in Teams
@@
Conflicts are inherent in all teams, whether human-only or human-AI. These conflicts typically arise from task, relational, and process-related dynamics (de Wit et al. 2013). In human-AI teams, such conflicts will continue to affect performance, but the generative or causal factors require new theoretical insights. Human-AI teams present new nuances, such as increased role fluidity—the ability to adapt and transition roles based on context and task requirements. Role fluidity can influence process clarity or conflict depending on how teams implement and manage their roles. If not managed early on, process conflict contributes to relational conflict. As AI agents take on more tasks, it is reasonable to expect changes to the shared mental models among team members. To the extent the mental models cohere, reductions in task conflicts are to be expected, which in turn may reduce process conflicts. It is, however, unclear as to whether the use of AI agents in teams interrupt the formation of shared mental models among humans or if they strengthen shared mental models. Below, we expand our discussion and outline key research directions, as summarized in Figure 4.
@@ 4.4.1   Expected Outcomes
Research on conflict in human-AI teams will need new theoretical lenses to study causative factors of conflict in such teams. The expected benefits of flat hierarchies and role fluidity have to be balanced against the potential harmful implications on team conflict. Extant research has raised the plausibility of increased interpersonal and status conflicts in flat organizational structures (Anicich et al. 2016). Fluidity in roles often causes process conflict, which, if not managed, leads to relational conflicts. If the key benefits of human-AI teams are to be realized, teams need to more effectively manage such conflicts. Human-AI teams need to address imbalances in responsibility and rewards to sustain trust, engagement, and overall team effectiveness (Gabriel et al. 2024). We argue that human-AI teams need their leaders and team members to develop a stronger meta-understanding of process goals and roles as opposed to focusing on transactional level agreements on process and task execution. This, in our opinion, is the key differentiation in theorizing the social dynamics of human-AI teams.
[…]
Figure 4: A Framework for Team Maintenance and Evolution in Human–AI Teams
[…]
@@ 4.4.2   Evaluations and Issues
Evaluating Conflict in Human-AI Teams:  Role fluidity enhances flexibility in task assignments and delegation but can lead to relational, process, and status-related conflicts if not well managed. Researchers should (1) explore how role fluidity leads to conflicts in human-AI teams and identify managerial actions that reduce such conflicts. For example, AI agents can aid shared mental model formation through self-explanation and reflection, improving role clarity and minimizing conflicts. Attitudes toward both human and AI teammates, especially when performance varies due to task allocations, are critical. For example, SimanTov-Nachlieli (2024) emphasize that employees’ trust in AI is linked to their performance rankings, making it vital to address dynamic perceptions of self-efficacy and collaboration willingness to maintain overall team resilience. Importantly, future research should also (2) evaluate how divergent and convergent thinking styles are balanced within human-AI teams to reduce task and process conflicts, focusing on the role of AI in fostering or disrupting this balance (Lanaj et al. 2018). This includes investigating how AI systems can facilitate social learning processes by enabling team members to integrate diverse perspectives, resolve conflicts constructively, and adapt roles dynamically for cohesive decision-making (Choi et al. 2024).
@@
Information Sharing and Knowledge Management in Human-AI Teams:   Effective information sharing is fundamental to human-AI team coordination and performance, particularly for enabling role fluidity and adaptive learning. Knowledge reuse involves not only accessing archived information but also meaningfully applying it to new tasks, allowing the team to build on past experiences and insights. We draw upon Nonaka’s dynamic theory of knowledge creation to identify key research directions for human-AI teams context (Nonaka 1994). Nonaka’s model includes four knowledge interaction patterns: socialization (tacit–tacit), externalization (tacit–explicit), internalization (explicit–tacit), and combination (explicit–explicit). The inclusion of AI agents clearly has ramifications on the interaction patterns in human-AI teams as it relates to knowledge sharing and creation. Of the four patterns, perhaps the combination pattern is relatively well understood. AI agents today are highly capable of combining internal and external knowledge and synthesizing it for human consumption. Future research should (1) examine how human-AI teams can effectively organize to improve socialization, externalization, and internalization. For example, can AI agents impede tacit knowledge formation in teams since they may interrupt or replace socialization between team members? how can organizations enhance socialization in AI-enabled teams? Additionally, (2) can the ability of AI agents to transform multimodal inputs (text, voice, video, etc.) accelerate knowledge creation through improved externalization? Similarly, (3) do interactions with AI agents during task completion impede or help internalization? Overall, human-AI teams will witness a disruption of the dyadic interactions in knowledge formation, and theoretical development in this domain is limited, making it difficult to reliably predict how organizational knowledge creation improves or deteriorates in the presence of AI agents.
@@ 4.4.3   Mechanisms and Resolutions
Fostering Adaptability and Resilience through Social Learning:  Resilient human-AI teams thrive on structured social learning processes such as reflection, experimentation, and collaborative problem-solving, which enable them to co-adapt and address disruptions effectively. Team reflexivity – collective reflection on objectives, strategies, and processes – enhances problem identification and resolution, improving overall team adaptability (Schippers et al. 2014). Collaborative problem-solving fosters critical thinking and collective learning, equipping teams to navigate dynamic and uncertain environments (Van den Bossche et al. 2006). These processes are integral to dynamically aligning roles between human and AI team members, strengthening team resilience. Particularly, simulated learning environments offer a practical and scalable strategy for accelerating these social learning processes. By providing safe, controlled spaces for experimentation, these environments enable teams to engage in iterative practice, refine workflows, and test strategies without the risks of real-world tasks (Salas et al. 2009). Simulation-based learning bridges theoretical knowledge and practical application, enhancing decision-making and adaptability in complex scenarios (Wei et al. 2024). Iterative experimentation within such environments fosters mutual adaptation between humans and AI, supporting the development of shared understanding and improved collaborative efficiency. Psychological safety is a cornerstone of effective social learning. Creating environments where team members can share honest feedback and take calculated risks without fear of negative repercussions promotes openness, creativity, and innovation (Edmondson 1999). Integrating simulated learning opportunities with feedback-rich environments ensures that human and AI agents can coevolve efficiently, addressing both operational efficiency and ethical alignment (Brinkmann et al. 2023). Future research should (1) investigate how simulated learning environments can incorporate iterative experimentation – such as real-time role rotation and collaborative problem-solving exercises – to enhance role fluidity and team resilience. Additionally, (2) explore governance strategies that embed social learning processes, such as feedback loops and ethical audits, into human-AI workflows to improve adaptability, trust, and long-term alignment (Mökander and Floridi 2023).
@@ 4.4.4   Discussion Extension: Considering Goal and Task Dynamics
On top of the discussions in Sections 4.4.2 and 4.4.3, future research should (1) explore how advanced social learning frameworks can integrate conflict management and iterative feedback loops into team maintenance and training, ensuring that human-AI teams maintain alignment and cohesion as goals and tasks evolve. Additionally, (2) investigate mechanisms to enhance role fluidity by developing adaptive protocols for task transitions, interdependency management, and real-time adjustments, ensuring that both human and AI contributions remain dynamically aligned with team objectives. These efforts will provide actionable insights into sustaining the effectiveness, adaptability, and resilience of human-AI teams in increasingly complex and dynamic environments.
@@ 5 Concluding Remarks
The disruptive innovation of AI agents compels us to perform a renewed examination of human–AI teaming. Through an extended team SA model, we present a comprehensive framework for examining key dimensions of human-AI teaming, including formulation, coordination, maintenance, and training. We also address issues arising from team complexity, such as variations in composition, size, goals, task dynamics, and long-term interactions. Ultimately, we emphasize the importance of revisiting human-AI teaming to foster a more dynamic and interactive relationship between humans and AI agents. Our work offers a forward-looking research agenda to guide future investigations in this evolving field.
Our discussion proposes the opportunities and challenges presented by the integration of AI systems into collaborative settings. On the one hand, AI offers immense potential for augmenting human capabilities, improving decision-making, and fostering innovation. On the other hand, unresolved tensions, such as trust and alignment considerations, underscore the need to design systems and processes that facilitate effective and sustainable collaboration and dynamic learning in teams. As such, future research should prioritize both technical development and the socio-behavioral foundations of teaming.
@@
In conclusion, human-AI teaming is a critical frontier in team science, offering vast opportunities for innovation and productivity while raising questions about the nature of collaboration, decision-making, and shared accountability. By addressing the gaps and tensions identified in this paper, researchers and practitioners can shape a future where human-AI teams achieve their full potential, benefiting individuals, organizations, and society at large.